Can you please help solve this?
Paula, Dianna and Annie are three friends. Paula is twice as old as Dianna, and in 2 years she will be three times as old as Annie. If Dianna is 5 years older than Annie, how old is Paula?
Thanks in advance

Found 2 solutions by Boreal, addingup:
Answer by Boreal(15235)

Dianna=x; Annie=A
Paula=2x
2x+2=3(A+2); remember, Annie is two years older in two years, too.
x-5=A
substitute
2x+2=3(x-5+2)=3x-9
x=11; Dianna is 11
Paula is 22
Annie is 6
In two years, Paula will be 24 and Annie 8, 3 times as old.

Answer by addingup(3677)

Paula: p
Diana: d
Annie: a
----------------------------
p = 2d or d = p/2
p+2 = 3(a+2)
d = a+5 or a = d-5
----------------------------
p+2 = 3(d-5+2)
add -5+2 and substitute for d:
p+2 = 3(p/2-3)
p+2 = 3/2p-9
1p-3/2p = -11
-p/2 = -11 multiply both sides times -1
p/2 = 11 NOTE: p/2 = Dianna's age
p = 22 Paula's age
Annie: 11-5 = 6
-------------------
Check:
In 2 years Paula will be 3 times as old as Annie:
p+2 = 3(a+2)
22+2 = 18+6 Correct, 24=24